# A Comprehensive Survey and Taxonomy on Point Cloud Registration Based on Deep Learning

### v1

@@ 3.3 Outlier Filtering
In PCR, outliers are defined as points lacking a corresponding counterpart. The principal objective of outlier filtering is to meticulously remove these outliers. Given their substantial influence on the outcomes of registration processes, the effective elimination of outliers is imperative to guarantee both robustness and accuracy. To identify outliers, 3DRegNet (Pais et al., 2020) utilizes a deep neural network to estimate the probability of a point being classified as an outlier, which effectively minimizes the influence of hypothetical outliers during the registration procedure. DHVR (Lee et al., 2021) places the initially predicted correspondences into a Hough voting module. This module casts votes in a deliberately sparse transformation parameter space, enhancing the accurate identification of inliers. Moreover, DLF (Zhang et al., 2022a) utilizes a classifier that combines the stacked order-aware modules to evaluate hypothesized outliers and determine the compatibility of hypothesized inliers.
The above methods directly estimate outliers after extracting features. However, during the feature extraction phase, they predominantly rely on methods like multilayer perceptron, inadvertently overlooking the critical aspect of the 3D spatial information. Furthermore, in classifying these features, each pair is assessed separately, ignoring the important consistency of inliers (Bai et al., 2021). Based on the above thinking, PointDSC (Bai et al., 2021) is proposed, which explicitly exploits the spatial compatibility inherently constructed by distance. It argues that not only should the relative distances of inliers between the point clouds to be registered remain consistent, but there also exists an inherent relationship among inliers within the single point cloud. Based on spatial compatibility, a second order spatial compatibility (Chen et al., 2022b) is proposed, which begins by converting the spatial compatibility matrix into a binary form and then calculates the similarity between two corresponding points based on the count of their mutually compatible points. This approach focusing on global rather than local compatibility, enhances early-stage differentiation between inliers and outliers. MAC (Zhang et al., 2023a) loosens the maximum clique constraint and mines more local consistency information in the compatibility graph for accurate pose hypothesis generation.
@@ 3.4 Transformation Parameter Estimation
The calculation of transformation parameters serves as the final step in PCR, with the widely adopted methods including RANSAC (Fischler and Bolles, 1981) and singular value decomposition (SVD) (Arun et al., 1987). RANSAC is commonly employed during the coarse registration stage to mitigate the impact of outliers, and it requires a predetermined number of iterations to solve. Unlike RANSAC, SVD does not necessitate an iterative solution. It estimates the transformation parameters directly based on the pose difference between the two point clouds, thus requiring a reliable feature extraction network for accurate results (Zhang et al., 2022b). The process of solving SVD reveals that the rotation matrix is computed prior to the calculation of the translation vector.
With the development of DL, some approaches strive to solve both rotation matrix and translation vector simultaneously using convolutional neural network (Deng et al., 2018; Pais et al., 2020). The effectiveness of this idea is examined across multiple models. However, simultaneous resolution of transformation parameters can lead to mutual interference (Chen et al., 2022c). To address this issue, DetarNet (Chen et al., 2022c) employs Siamese networks to independently decouple transformation parameters in a two-step process. Initially, a regression network computes the translation vector, followed by the utilization of SVD to determine the rotation matrix. FINet (Xu et al., 2022) leverages point-wise and global features to enhance information association between point clouds to be registered at multiple stages. At the same time, a dual-branch structure containing a rotation regression branch and a translation regression branch is designed to predict the rotation matrix and translation vector, respectively.
@@ 3.5.1 ICP-based
While the classical ICP algorithm may be less effective than DL-based algorithms for PCR tasks, it still offers valuable advantages worth exploring. One such advantage is its iterative optimization thought, which has been widely adopted in various methods to refine estimation transformation parameters. In DCP (Wang and Solomon, 2019a) and PRNet (Wang and Solomon, 2019b), the entire network follows an iterative process to enhance the initial prediction of the rotation matrix and translation vector, progressively refining them from a coarse to a fine level. More specifically, before each iteration, the point cloud to be registered is updated with the transformation parameters estimated in the previous iteration. This incremental refinement process allows for the gradual improvement of the predicted transformation parameters. Contrasting with the previously discussed algorithms that iterate through the entire network, IDAM (Li et al., 2020a) distinctively positions the feature extraction component outside the iterative loop, which reduces the computational burden to a certain extent. Furthermore, it integrates distance information into the iterative network and incorporates a two-stage point elimination module. This design effectively filters out points that are detrimental to the registration process.
3.5.2 Probabilistic-based
@@
Probability-based PCR algorithms integrate probabilistic knowledge within the registration framework to enhance the optimization process. These algorithms typically utilize probabilistic models to depict the matching relationship and inherent uncertainty between point clouds to be registered.
@@
As a commonly used probability model, the Gaussian mixture model (GMM) finds optimal alignments by integrating an expectation-maximization (EM) method into a maximum likelihood framework (Eckart et al., 2018). However, the EM process can be computationally intensive and potentially lead to incorrect data associations, especially in registrations with significant angular disparities (Yuan et al., 2020). To address the aforementioned challenges, a technique called deep Gaussian mixture registration (DeepGMR) (Yuan et al., 2020) is proposed, which leverages a neural network to search correspondences between points and GMM parameters. Furthermore, two differentiable modules are employed to estimate the optimal transformation parameters. OGMM (Mei et al., 2023a) utilizes predictions of the overlapping area between two input point clouds for GMM estimation, framing the registration task as minimizing the variance between the two GMMs. In (Chen et al., 2023a), GMM is formulated as a distribution that encompasses comprehensive representation capabilities, incorporating both global and local information.
@@
In addition to GMM, the Bayesian probabilistic model is also utilized for PCR. VBReg (Jiang et al., 2023) introduces a variable non-local network architecture, which employs variational Bayesian inference for non-local feature learning. This approach enables the modeling of Bayesian-driven long-range dependencies and facilitates the acquisition of discriminative feature representations for inlier/outlier.
@@ 3.6 Multimodal
PCR-CG (Zhang et al., 2022c) and PEAL (Yu et al., 2023b) employ a two-dimensional (2D) image matching technique to establish 2D correspondences, which are then projected onto point clouds using a 2D to 3D projection module, facilitating the identification of overlapping regions. ImLoveNet (Chen et al., 2022a) also utilizes images to enhance predictions in overlapping regions, directly employing cross-fusion technology to amalgamate the 3D features extracted directly from point clouds with the 3D features simulated from two-dimensional features derived from images. IMFNET (Huang et al., 2022c) proposes an interpretable module to explain the contribution of the original points to the final descriptor. This approach significantly enhances both the transparency and effectiveness of the descriptor. GMF (Huang et al., 2022b) integrates texture and structural information through a cross-attention fusion layer. Additionally, it incorporates a convolutional position encoding layer, which is instrumental in accentuating distinctions and focusing on neighboring information. Consequently, these enhancements contribute to improving correspondence quality and standard accuracy in the model.
@@ 4.1 Correspondence-free
In general, correspondence-based unsupervised registration methods first extract global features from the source and target point clouds and then minimize the difference between them to regress the transformation parameters. As depicted in Figure 3, the correspondence-free algorithms utilize the calculation of the distance between point clouds to define the loss function, termed distance loss. Typically, distance loss in the unsupervised methods uses CD, which measures the distance or feature differences between point pairs in two point clouds bidirectionally, with its formula is
@@
In the field of correspondence-free unsupervised methods, an early significant contribution is PPF-FoldNet (Deng et al., 2018), which starts by constructing four-dimensional (4D) point-pair features. These features are subsequently fed into an end-to-end architecture resembling a folded network, utilizing an encoder-decoder structure for reconstruction. The loss function involves comparing the CD between the 4D point pair features before and after reconstruction. Sun et al. (Sun et al., 2023) have further developed the PointNetLK algorithm for use in cross-source PCR, employing global features for CD calculation. UGMM (Huang et al., 2022a) presents a novel approach, redefining the PCR challenge as a clustering problem and estimating posterior probabilities through unsupervised learning. This method uses the CD between Gaussian mixtures derived from the point clouds as the loss function. UPCR (Zhang et al., 2021) introduces dual point cloud representations: pose-invariant and pose-related. The pose-related representations are leveraged to learn relative poses, which are essential for deriving transformation parameters. Moreover, the CD is also integrated into the loss function to evaluate the discrepancy between the source point cloud and the target point cloud. PCRNet (Sarode et al., 2019), while following a similar approach as UPCR in designing its loss function, distinguishes itself by utilizing the earth mover’s distance.
@@ 4.2 Correspondence-based
Compared with the correspondence-free unsupervised method, the correspondence-based unsupervised method first extracts features, and then uses the correspondence relative step in Figure 3 (including correspondence search or outlier filtering) to establish point-level, distribution-level, or cluster-level correspondences. Finally, the rigid transformation parameters are estimated from these correspondences. CEMNet (Jiang et al., 2021) integrates the scaling estimator into the function that measures the registration error to weaken the negative impact of outliers on registration accuracy. CEMNet also uses CD as the loss function.
In addition to CD, correspondence-based unsupervised algorithms also designed various other losses to refine aligned point clouds. RIENet (Shen et al., 2022) proposes a reliable inlier estimation module and designs the neighborhood consensus loss and spatial consistency loss to reduce the local differences and global differences of the point cloud to be registered. UDPReg (Mei et al., 2023b) finds correspondences from cluster-level and point-level, and designs self-consistency loss, cross-consistency loss, and local contrastive loss to enable unsupervised learning.
@@ 5 Challenges and Opportunities
  • Towards realistic data generation: A major challenge is bridging the gap between synthetic and real-world data. Most methods often rely on Gaussian noise to mimic realistic data, which fails to capture the complexity of actual data. Chen et al. (Chen et al., 2023c) propose a new perspective that introduces the diffusion model to generate noisy training data. Future research can focus on integrating other generative models to simulate noise and occlusions, or developing data generation methodologies that can simulate realistic data independently of external networks.
@@
  • Designing new metrics: (Chen et al., 2023d) designed a new metric that effectively achieves dual optimization in processing speed and registration accuracy. This advancement not only enhances the performance of existing registration networks but also opens new perspectives for PCR tasks. Future research can explore innovative evaluation metrics that comprehensively consider factors such as runtime speed, model size, and registration quality.
@@ Acknowledgments
This work was supported in part by the grant of the National Science Foundation of China under Grant 62172090.

### v2

@@ 3.3 Outlier Filtering
In PCR, outliers are defined as points lacking a corresponding counterpart. The principal objective of outlier filtering is to meticulously remove these outliers. Given their substantial influence on the outcomes of registration processes, the effective elimination of outliers is imperative to guarantee both robustness and accuracy. To identify outliers, 3DRegNet Pais et al. (2020) utilizes a deep neural network to estimate the probability of a point being classified as an outlier, which effectively minimizes the influence of hypothetical outliers during the registration procedure. DHVR Lee et al. (2021) places the initially predicted correspondences into a Hough voting module. This module casts votes in a deliberately sparse transformation parameter space, enhancing the accurate identification of inliers. Moreover, DLF Zhang et al. (2022a) utilizes a classifier that combines the stacked order-aware modules to evaluate hypothesized outliers and determine the compatibility of hypothesized inliers.
The above methods directly estimate outliers after extracting features. However, during the feature extraction phase, they predominantly rely on methods like multilayer perceptron, inadvertently overlooking the critical aspect of the 3D spatial information. Furthermore, in classifying these features, each pair is assessed separately, ignoring the important consistency of inliers Bai et al. (2021). Based on the above thinking, PointDSC Bai et al. (2021) is proposed, which explicitly exploits the spatial compatibility inherently constructed by distance. It argues that not only should the relative distances of inliers between the point clouds to be registered remain consistent, but there also exists an inherent relationship among inliers within the single point cloud. Based on spatial compatibility, a second order spatial compatibility Chen et al. (2022b) is proposed, which begins by converting the spatial compatibility matrix into a binary form and then calculates the similarity between two corresponding points based on the count of their mutually compatible points. This approach focusing on global rather than local compatibility, enhances early-stage differentiation between inliers and outliers. MAC Zhang et al. (2023a) loosens the maximum clique constraint and mines more local consistency information in the compatibility graph for accurate pose hypothesis generation.
@@ 3.4 Transformation Parameter Estimation
The calculation of transformation parameters serves as the final step in PCR, with the widely adopted methods including RANSAC Fischler and Bolles (1981) and singular value decomposition (SVD) Arun et al. (1987). RANSAC is commonly employed during the coarse registration stage to mitigate the impact of outliers, and it requires a predetermined number of iterations to solve. Unlike RANSAC, SVD does not necessitate an iterative solution. It estimates the transformation parameters directly based on the pose difference between the two point clouds, thus requiring a reliable feature extraction network for accurate results Zhang et al. (2022b). The process of solving SVD reveals that the rotation matrix is computed prior to the calculation of the translation vector.
With the development of DL, some approaches strive to solve both rotation matrix and translation vector simultaneously using convolutional neural network Deng et al. (2018); Pais et al. (2020). The effectiveness of this idea is examined across multiple models. However, simultaneous resolution of transformation parameters can lead to mutual interference Chen et al. (2022c). To address this issue, DetarNet Chen et al. (2022c) employs Siamese networks to independently decouple transformation parameters in a two-step process. Initially, a regression network computes the translation vector, followed by the utilization of SVD to determine the rotation matrix. FINet Xu et al. (2022) leverages point-wise and global features to enhance information association between point clouds to be registered at multiple stages. At the same time, a dual-branch structure containing a rotation regression branch and a translation regression branch is designed to predict the rotation matrix and translation vector, respectively.
@@ 3.5 Optimization
ICP-based. While the classical ICP algorithm may be less effective than DL-based algorithms for PCR tasks, it still offers valuable advantages worth exploring. One such advantage is its iterative optimization thought, which has been widely adopted in various methods to refine estimation transformation parameters. In DCP Wang and Solomon (2019a) and PRNet Wang and Solomon (2019b), the entire network follows an iterative process to enhance the initial prediction of the rotation matrix and translation vector, progressively refining them from a coarse to a fine level. More specifically, before each iteration, the point cloud to be registered is updated with the transformation parameters estimated in the previous iteration. This incremental refinement process allows for the gradual improvement of the predicted transformation parameters. Contrasting with the previously discussed algorithms that iterate through the entire network, IDAM Li et al. (2020a) distinctively positions the feature extraction component outside the iterative loop, which reduces the computational burden to a certain extent. Furthermore, it integrates distance information into the iterative network and incorporates a two-stage point elimination module. This design effectively filters out points that are detrimental to the registration process.
@@
Probabilistic-based. Probability-based PCR algorithms integrate probabilistic knowledge within the registration framework to enhance the optimization process. These algorithms typically utilize probabilistic models to depict the matching relationship and inherent uncertainty between point clouds to be registered.
@@
As a commonly used probability model, the Gaussian mixture model (GMM) finds optimal alignments by integrating an expectation-maximization (EM) method into a maximum likelihood framework Eckart et al. (2018). However, the EM process can be computationally intensive and potentially lead to incorrect data associations, especially in registrations with significant angular disparities Yuan et al. (2020). To address the aforementioned challenges, a technique called deep Gaussian mixture registration (DeepGMR) Yuan et al. (2020) is proposed, which leverages a neural network to search correspondences between points and GMM parameters. Furthermore, two differentiable modules are employed to estimate the optimal transformation parameters. OGMM Mei et al. (2023a) utilizes predictions of the overlapping area between two input point clouds for GMM estimation, framing the registration task as minimizing the variance between the two GMMs. In Chen et al. (2023a), GMM is formulated as a distribution that encompasses comprehensive representation capabilities, incorporating both global and local information.
@@
In addition to GMM, the Bayesian probabilistic model is also utilized for PCR. VBReg Jiang et al. (2023) introduces a variable non-local network architecture, which employs variational Bayesian inference for non-local feature learning. This approach enables the modeling of Bayesian-driven long-range dependencies and facilitates the acquisition of discriminative feature representations for inlier/outlier.
@@ 3.6 Multimodal
PCR-CG Zhang et al. (2022c) and PEAL Yu et al. (2023b) employ a two-dimensional (2D) image matching technique to establish 2D correspondences, which are then projected onto point clouds using a 2D to 3D projection module, facilitating the identification of overlapping regions. ImLoveNet Chen et al. (2022a) also utilizes images to enhance predictions in overlapping regions, directly employing cross-fusion technology to amalgamate the 3D features extracted directly from point clouds with the 3D features simulated from two-dimensional features derived from images. IMFNET Huang et al. (2022c) proposes an interpretable module to explain the contribution of the original points to the final descriptor. This approach significantly enhances both the transparency and effectiveness of the descriptor. GMF Huang et al. (2022b) integrates texture and structural information through a cross-attention fusion layer. Additionally, it incorporates a convolutional position encoding layer, which is instrumental in accentuating distinctions and focusing on neighboring information. Consequently, these enhancements contribute to improving correspondence quality and standard accuracy in the model.
@@ 4.1 Correspondence-free
In general, correspondence-based unsupervised methods first extract global features from the source and target point clouds and then minimize the difference between them to regress the transformation parameters. As depicted in Figure 3, these algorithms utilize the calculation of the distance between point clouds to define the loss function, termed distance loss. Typically, distance loss in the unsupervised methods uses CD, which measures the distance or feature differences between point pairs in two point clouds bidirectionally, formulated as
@@
In the field of correspondence-free unsupervised methods, an early significant contribution is PPF-FoldNet Deng et al. (2018), which starts by constructing four-dimensional (4D) point-pair features. These features are subsequently fed into an end-to-end architecture resembling a folded network, utilizing an encoder-decoder structure for reconstruction. The loss function involves comparing the CD between the 4D point pair features before and after reconstruction. Sun et al. Sun et al. (2023) have further developed the PointNetLK algorithm for use in cross-source PCR, employing global features for CD calculation. UGMM Huang et al. (2022a) presents a novel approach, redefining the PCR challenge as a clustering problem and estimating posterior probabilities through unsupervised learning. This method uses the CD between Gaussian mixtures derived from the point clouds as the loss function. UPCR Zhang et al. (2021) introduces dual point cloud representations: pose-invariant and pose-related. The pose-related representations are leveraged to learn relative poses, which are essential for deriving transformation parameters. Moreover, the CD is also integrated into the loss function to evaluate the discrepancy between the source point cloud and the target point cloud. PCRNet Sarode et al. (2019), while following a similar approach as UPCR in designing its loss function, distinguishes itself by utilizing the earth mover’s distance.
@@ 4.2 Correspondence-based
Compared with the correspondence-free unsupervised method, the correspondence-based unsupervised method first extracts features, and then uses the correspondence relative step in Figure 3 (including correspondence search or outlier filtering) to establish point-level, distribution-level, or cluster-level correspondences. Finally, the rigid transformation parameters are estimated from these correspondences. CEMNet Jiang et al. (2021) integrates the scaling estimator into the function that measures the registration error to weaken the negative impact of outliers on registration accuracy. CEMNet also uses CD as the loss function.
In addition to CD, correspondence-based unsupervised algorithms also designed various other losses to refine aligned point clouds. RIENet Shen et al. (2022) proposes a reliable inlier estimation module and designs the neighborhood consensus loss and spatial consistency loss to reduce the local differences and global differences of the point cloud to be registered. UDPReg Mei et al. (2023b) finds correspondences from cluster-level and point-level, and designs self-consistency loss, cross-consistency loss, and local contrastive loss to enable unsupervised learning.
@@ 5 Challenges and Opportunities
  • Towards realistic data generation: A major challenge is bridging the gap between synthetic and real-world data. Most methods often rely on Gaussian noise to mimic realistic data, which fails to capture the complexity of actual data. Chen et al. Chen et al. (2023c) propose a new perspective that introduces the diffusion model to generate noisy data. Future research can focus on integrating other generative models to simulate noise and occlusions, or developing data generation methods that can simulate realistic data independently of external networks.
@@
  • Designing new metrics: Chen et al. (2023d) designed a new metric that effectively achieves dual optimization in processing speed and registration accuracy. This advancement not only enhances the performance of existing registration networks but also opens new perspectives for PCR tasks. Future research can explore innovative evaluation metrics that comprehensively consider factors such as runtime speed, model size, and registration quality.
@@ Acknowledgments
This work was supported in part by the grant of the National Science Foundation of China under Grant 62172090. We thank the Big Data Computing Center of Southeast University for providing the facility support in this paper.